OpenStudy (anonymous):

B is between A and C. Find BC if AB = 3, BC = 4x + 1, AC = 8. 1) 1 2) 5 3) -5 4) -1

OpenStudy (aroub):

ab+bc=ac 3+4x+1=8 4+4x=8 4x=8-4 4x=4 x=1

OpenStudy (aroub):

so its 1

OpenStudy (anonymous):

thank you!

OpenStudy (aroub):

anytime :)